# Benchmarking General-Purpose In-Context Learning

### v5

@@ A.2 Data Collection Strategy for Maze World
For dataset collection and static evaluation of Maze World, inspired by dataset aggregation [Ross et al., 2011] and noise distillation [Zisman et al., 2023], we utilize distinct behavior and reference policies. The behavior policy is determined by previleged agent where $p(\text{STM}\rightarrow\text{LTM})$ randomly sampled between $[0,50\%]$. Additionally, with the probability of $\epsilon\in[0,80\%]$ the behavior action is determined by random decision. With the states generated by behavior policy, we use the reference policy with $p(\text{STM}\rightarrow\text{LTM})=100\%$ to label the action.
@@ A.5 Additional Diversity in the Maze World
It is worth noting that Maze World is a lightweight and fast engine with an easy-to-use Python API based on GYM [Brockman et al., 2016] for simulating navigation in various mazes. It runs smoothly at over 20 frames per second on a standard CPU. In addition to the aforementioned details, Maze World incorporates various configurations and hyper-parameters that can be adjusted to create a wide range of tasks. For example, it offers different types of observation spaces and action spaces, including discrete and continuous actions, as well as 2D and 3D observations (Figure 16a). Beyond navigation tasks, we introduce more challenging survival tasks, where the rewards provided by the PNTs are static within an episode but initially unknown (can be negative, Figure 16b). The basic environments can also be modified in terms of the size of the basic cell (Figure 16c,16d), the height of the ceiling (Figure 16e,16f), the density of obstacles (Figure 16g,16h), and the range of view (Figure 16i,16j).
@@ A.6 Details of Task Configurations
In Table 2, we list the detailed settings for the tasks used in this paper. The rewards for reaching the target are by default related to the size of the maze. Additionally, we note that the density of obstacles controls the number of loops in the environments. We recommend using a density between 0.15 and 0.40. A density over 0.50 results in mazes without any loops, significantly reducing the difficulty of traversing the maze (or performing exploration), as an agent can easily traverse the entire world by always going right (or left). We believe this might be a crucial reason why, in Morad et al. [2023], where the agent is only required to navigate to a single target, context-free models can outperform memory-augmented models.

### v6

@@ A.2 Data Collection Strategy for Maze World
For dataset collection and static evaluation of Maze World, inspired by dataset aggregation [37] and noise distillation [55], we utilize distinct behavior and reference policies. The behavior policy is determined by previleged agent where $p(\text{STM}\rightarrow\text{LTM})$ randomly sampled between $[0,50\%]$. Additionally, with the probability of $\epsilon\in[0,80\%]$ the behavior action is determined by random decision. With the states generated by behavior policy, we use the reference policy with $p(\text{STM}\rightarrow\text{LTM})=100\%$ to label the action.
@@ A.5 Additional Diversity in the Maze World
It is worth noting that Maze World is a lightweight and fast engine with an easy-to-use Python API based on GYM [2] for simulating navigation in various mazes. It runs smoothly at over 20 frames per second on a standard CPU. In addition to the aforementioned details, Maze World incorporates various configurations and hyper-parameters that can be adjusted to create a wide range of tasks. For example, it offers different types of observation spaces and action spaces, including discrete and continuous actions, as well as 2D and 3D observations (Figure 19a). Beyond navigation tasks, we introduce more challenging survival tasks, where the rewards provided by the PNTs are static within an episode but initially unknown (can be negative, Figure 19b). The basic environments can also be modified in terms of the size of the basic cell (Figure 19c,19d), the height of the ceiling (Figure 19e,19f), the density of obstacles (Figure 19g,19h), and the range of view (Figure 19i,19j).
@@ A.6 Details of Task Configurations
In Table 2, we list the detailed settings for the tasks used in this paper. The rewards for reaching the target are by default related to the size of the maze. Additionally, we note that the density of obstacles controls the number of loops in the environments. We recommend using a density between 0.15 and 0.40. A density over 0.50 results in mazes without any loops, significantly reducing the difficulty of traversing the maze (or performing exploration), as an agent can easily traverse the entire world by always going right (or left). We believe this might be a crucial reason why, in [29], where the agent is only required to navigate to a single target, context-free models can outperform memory-augmented models.